OpenStudy (anonymous):

9-3(7x-4)+10 1/3x=-24 can anyone show me the steps to this multi step equation

OpenStudy (anonymous):

\[9 - 3(7x - 4) + 10\frac{ 1 }{ 3x } = -24\] ^^^^^^ that?

OpenStudy (anonymous):

9-3(7x-4)+(10 1/3)x=-24

OpenStudy (anonymous):

is it \[10\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

yes x is at the end

OpenStudy (anonymous):

\[9 - 3(7x - 4) + (10\frac{ 1 }{ 3 })x = -24\]

OpenStudy (anonymous):

yes just like that , but technically there are no () around the 10 1/3 x

OpenStudy (anonymous):

so which is it!

OpenStudy (anonymous):

is it my first post?

OpenStudy (anonymous):

Are you saying that its 9-3(7x-4) +10 1/3=-24 but in your book or on your paper the () isnt around the "10 1/3"?

OpenStudy (anonymous):

\[10{ }{ \frac{ 1}{ 3} }x\]

OpenStudy (anonymous):

yes no () around the fraction